Two charges are located on the positive x-axis of a coordinate system. Charge π‘ž1 = 2 Γ— 10βˆ’9𝐢 is 2cm from the origin, and charge π‘ž2 = βˆ’3 Γ— 10βˆ’9𝐢 is 4cm from the origin. What is the total force exerted by these two charges on a charge π‘ž3 = 5 Γ— 10βˆ’9𝐢 located at the origin?


1
Expert's answer
2022-03-15T10:34:30-0400

Solution.

"q_1=2\sdot10^{-9}C;
r_1=0.02m;"

"q_2=-3\sdot10^{-9}C; r_2=0.04m;"

"F_{13}=k \dfrac{q_1\sdot q_3}{r_1^2};"

"F_{13}=9\sdot10^9\dfrac{2\sdot10^{-9}\sdot5\sdot10^{-9}}{0.02^2}=2.25\sdot10^{-4}N;"

"F_{23}=k\dfrac{q_2\sdot q_3}{r_2^2};"

"F_{23}=9\sdot10^9\dfrac{3\sdot10^{-9}\sdot5\sdot10^{-9}}{0.04^2}=0.84\sdot10^{-4}N;"

"F=F_{13}-F_{23}=2.25\sdot 10^{-9} - 0.84\sdot 10^{-9}=1.41\sdot10^{-9}N;"

Answer: "F=1.41\sdot10^{-9}N."